OpenStudy (anonymous):

Two blocks of mass M and 3M are connected by a light chord which passes over a light frictionless pully. the blocks are released at the same height at t=0. What is the magnitude of acceleration of the centre of mass of the two block system? I got the answer as g/4 but the answerkey show as g/2. Which is correct? Please explain

OpenStudy (aravindg):

Could you show your attempt with the steps included? I can debug it for you. Please also show the free body diagrams.

OpenStudy (anonymous):

Dear AravindG, My attempt was :- Let T be tension and a be acceleration, then, 3Mg - T = 3Ma .....(1) T - Mg = Ma .....(2) adding 1&2 above, 2Mg = 4Ma a = g/2 ( acceleration of each block) Net downward force = 3Ma - Ma = 2Ma therefore, acceleration of center of mass = 2Ma/4M = a/2 = g/4 Please do correct me, Thanks!

OpenStudy (michele_laino):

hint: here are the equations of motion: \[\huge \left\{ \begin{gathered} \tau - Mg = Ma \hfill \\ \tau - 3Mg = - 3Ma \hfill \\ \end{gathered} \right.\]  please, solve such system for acceleration \(a\)